The weights of a women population are normally distributed with mean 60kg and standard deviation 5kg. Find the probability that the weight of a woman chosen at random is between 61 and 65kg.
1
Expert's answer
2020-10-20T18:54:07-0400

"\mu = 60"

"\sigma = 5"

"P(61<X<65)=P( \frac{61-\mu}{\sigma}<Z<\frac{65-\mu}{\sigma})=P( \frac{61-60}{5}<Z<\frac{65-60}{5})=P(0.2<Z<1) = P(Z<1) -P(Z<0.2)=0.8413-0.5793=0.262"

Answer: the probability that the weight of a woman chosen at random is between 61 and 65kg is 26.2%